OpenStudy (anonymous):

Find the difference. 14x - (x2 + 8x) =

OpenStudy (anonymous):

You mean 2x, right? If that's the case, then it's 14x-10x=4x.

OpenStudy (anonymous):

=6x-x^2=x(6-x)

OpenStudy (anonymous):

Oh, sorry, that's what he meant.

OpenStudy (eyust707):

this is one of those ones you just need to know the rules of algebra. 

OpenStudy (eyust707):

this negative sign right here is really like a -1